William sunk 75% of the shots he made on the basket. If he threw 160 shots, how many did he miss?
1
Expert's answer
2019-03-26T11:14:13-0400

Let X be the number of successes in n independent repetitions of a random experiment have probability of success = 0.75


"p={X \over n}"

Let Y be the number of misses in n independent repetitions. Then


"X+Y=n"

"Y=n-X=n-np=n(1-p)"

If William threw 160 shots, he did , how many did he miss?


"Y=160(1-0.75)=40"

William did 40 misses.